# Derived deformation functors, Koszul duality, and Maurer–Cartan spaces

### v1

@@ 5.2.2. Key features of the proof
That the adjunction $\mathrm{Tot}\,\dashv\mathrm{Tot}\,_{*}$ gives an $\infty$-equivalence is a more unusual argument [Pri07, Theorem 4.48]. The left Quillen functor $\mathrm{Tot}\,$ preserves fibrant objects, since they are cofree, and satisfies $\tan(\mathrm{Tot}\,C)\cong\mathrm{Tot}\,\tan(C)$. Moreover, $\mathrm{Tot}\,_{*}$ automatically preserves fibrant objects and has $\mathrm{H}_{i}\mathrm{H}^{0}\tan(\mathrm{Tot}\,_{*}B)\cong\mathrm{H}^{-i}B$ and $\mathrm{H}_{0}\mathrm{H}^{n}\tan(\mathrm{Tot}\,_{*}B)\cong\mathrm{H}^{n}B$ for such objects. For any fibrant $C\in\mathrm{dg_{+}DG^{+}\mathcal{C}C}^{\mathrm{nilp}}$ (resp. fibrant $B\in\mathrm{DG\mathcal{C}C}^{\mathrm{nilp}}$), the unit $C\to\mathrm{Tot}\,_{*}\mathrm{Tot}\,C$ (resp. co-unit $\mathrm{Tot}\,\mathrm{Tot}\,_{*}B\to B$) is then a weak equivalence between fibrant objects because it induces a $\mathrm{Tot}\,$-quasi-isomorphism (resp. quasi-isomorphism) on indecomposables.
@@ 5.3. Representability
Again, this step holds for any dg co-operad $\mathcal{C}$ concentrated in non-negative chain degrees, though it is also true much more generally. Analogues are given in [Pri07] for simplicial Artinian rings in finite and mixed characteristic, but the argument will apply to much more general Artinian categories with well-behaved analogues of the classes of central and acyclic central extensions.

### v2

@@ 5.2.2. Key features of the proof
That the adjunction $\mathrm{Tot}\,\dashv\mathrm{Tot}\,_{*}$ gives an $\infty$-equivalence is a more unusual argument [Pri07, Theorem 4.48]. The left Quillen functor $\mathrm{Tot}\,$ preserves fibrant objects, since they are cofree, and satisfies $\tan(\mathrm{Tot}\,C)\cong\mathrm{Tot}\,\tan(C)$. Moreover, $\mathrm{Tot}\,_{*}$ automatically preserves fibrant objects and has $\mathrm{H}_{i}\mathrm{H}^{0}\tan(\mathrm{Tot}\,_{*}B)\cong\mathrm{H}^{-i}\tan B$ and $\mathrm{H}_{0}\mathrm{H}^{n}\tan(\mathrm{Tot}\,_{*}B)\cong\mathrm{H}^{n}\tan B$ for such objects. For any fibrant $C\in\mathrm{dg_{+}DG^{+}\mathcal{C}C}^{\mathrm{nilp}}$ (resp. fibrant $B\in\mathrm{DG\mathcal{C}C}^{\mathrm{nilp}}$), the unit $C\to\mathrm{Tot}\,_{*}\mathrm{Tot}\,C$ (resp. co-unit $\mathrm{Tot}\,\mathrm{Tot}\,_{*}B\to B$) is then a weak equivalence between fibrant objects because it induces a $\mathrm{Tot}\,$-quasi-isomorphism (resp. quasi-isomorphism) on indecomposables.
@@ 5.3. Representability
Again, this step holds for any dg co-operad $\mathcal{C}$ concentrated in non-negative chain degrees, though it is also true much more generally. Analogues are given in [Pri07, Theorems 2.14 and 2.30] for simplicial Artinian rings in finite and mixed characteristic, but the argument should apply to much more general Artinian categories with well-behaved analogues of the classes of central and acyclic central extensions.
@@
Acknowledgements
I would like to thank the anonymous referees for diligently identifying omissions, oversights and particularly gaps in some arguments for generalising results to the stated generality.